### v2

# Radar Signal Recognition through Self-Supervised Learning and Domain Adaptation

Abstract
Automatic radar signal recognition (RSR) plays a pivotal role in electronic warfare (EW), as accurately classifying radar signals is critical for informing decision-making processes. Recent advances in deep learning have shown significant potential in improving RSR performance in domains with ample annotated data. However, these methods fall short in EW scenarios where annotated RF data are scarce or impractical to obtain. To address these challenges, we introduce a self-supervised learning (SSL) method which utilises masked signal modelling and RF domain adaption to enhance RSR performance in environments with limited RF samples and labels. Specifically, we investigate pre-training masked autoencoders (MAE) on baseband in-phase and quadrature (I/Q) signals from various RF domains and subsequently transfer the learned representation to the radar domain, where annotated data are limited. Empirical results show that our lightweight self-supervised ResNet model with domain adaptation achieves up to a 17.5% improvement in 1-shot classification accuracy when pre-trained on in-domain signals (i.e., radar signals) and up to a 16.31% improvement when pre-trained on out-of-domain signals (i.e., comm signals), compared to its baseline without SSL. We also provide reference results for several MAE designs and pre-training strategies, establishing a new benchmark for few-shot radar signal classification.
Index Terms:
few-shot, radar signal recognition, domain adaptation, self-supervised learning, masked autoencoder
I Introduction
Automatic radar signal recognition (RSR) is a crucial capability in cognitive electronic warfare (CEW) [1, 2], where accurate radar signal classification is essential for informed decision-making in the battlefield. Recent progress in deep learning has demonstrated significant potential [3] in addressing RSR sub-tasks, such as automatic modulation classification (AMC) [4] and radar activity segmentation [5], when abundant data is available for algorithm development. However, most existing RSR methods rely heavily on annotated data, which presents a challenge in EW scenarios where mission-specific radio frequency (RF) data is often scarce or difficult to acquire. Moreover, practical EW operations often demand rapid in-mission re-training of models on limited data to respond to the rapidly changing threat landscape [2]. This makes few-shot learning a critical yet underexplored area in RSR.
Figure 1: The proposed self-supervised learning architecture. Encoders are pre-trained using different masking strategies and masking ratios on various RF domains. Fine-tuning is conducted using a linear probing classifier.
In this paper, we introduce a self-supervised learning (SSL) approach that leverages masked signal modelling (MSM) and domain adaptation for few-shot RSR. Our proposed method pre-trains models on baseband I/Q signals with diverse characteristics from various RF domains. We then apply few-shot transfer learning to adapt pre-trained models to the radar domain using limited training samples. Our contributions are as follows: (i) we propose a modular and effective SSL architecture using different MSM strategies to enhance radar signal classification performance in few-shot settings; (ii) we introduce RF domain adaptation for self-supervised pre-training, enabling effective adaptation to RF domains with limited data; and (iii) we publish our SSL datasets [^1] Our datasets used for self-supervised pre-training, fine-tuning, and evaluation can be accessed at: https://github.com/abcxyzi/RadCharSSL accompanied by a new benchmark for few-shot radar signal classification.
II Related Work
Learning-based approaches for RF signal recognition have been an active research area over the past decade [4, 6, 7, 8, 9, 10, 11, 12]. Previous methods have relied on 2D feature transformations [13, 14, 15] to enhance classification performance in supervised settings. More recently, 1D approaches [10, 11, 16] have shown improved performance by effectively capturing fine-grained temporal relationships, as demonstrated in various signal recognition tasks within the RF domain, such as modulation classification [17, 9], radar signal characterisation [11], RF fingerprinting [18], and radar activity segmentation [5]. While unsupervised methods [19, 20] have been applied to AMC, self-supervised approaches have been more extensively studied in related signal processing domains, such as visual recognition [21, 22], audio recognition [23, 24], bio-signal classification [25, 26], and natural language processing [27]. Notably, masked autoencoders (MAE) have proven to be flexible and effective SSL architectures for time series classification [28] and representation learning in related tasks [29, 25, 24, 30]. Due to their success, MAE-based approaches have been adapted in recent work [31, 32] for RF emitter identification, particularly when training samples are limited. However, self-supervised approaches for RSR remain largely unexplored, likely due to the scarcity of publicly available radar datasets [3, 11]. Moreover, existing work in RSR has yet to explore SSL with RF domain adaptation.
III Proposed Method
III-A Autoencoder for I/Q Signals
Our SSL framework comprises two stages (Fig. 1). First, self-supervised pre-training is conducted via a masked autoencoder trained on a source domain. Then, the pre-trained encoder is fine-tuned on the target domain. We follow the modular encoder-decoder paradigm [33] to construct our autoencoder (Fig. 1). In the pre-training stage, we utilise asymmetric masked autoencoding [22] whereby the model operates on a partially observed I/Q signal. We train the model to reconstruct the original signal using a sample-wise similarity loss function [5]. This process is considered self-supervised as no annotations are required for the reconstruction task. After pre-training, we replace the decoder with a linear probing classifier consisting of a flatten operation followed by a fully connected layer. We fine-tune the classifier on the target domain with a small amount (i.e., few-shot) of annotated training data before evaluating it on a large test set. We implement several autoencoders to examine their effectiveness in self-supervised RSR. We focus only on lightweight models, as rapid retraining and processing speed are critical considerations for EW [2]. We adopt the 1D ResNet [31], the $2$-stage MS-TCN [34, 5], and the WaveNet [35, 36] as our baseline autoencoders. Pre-training of MS-TCN and WaveNet is conducted end-to-end as they do not have independent encoders and decoders. As such, the linear probing classifier is appended to these models during fine-tuning. In our experiments, we use the $\ell_{1}$ regression loss for pre-training and cross entropy loss for fine-tuning.
III-B Masked Signal Modelling
Masked signal modelling (MSM) is conceptually similar to masked image modelling (MIM) [37, 29], which has gained popularity as a simple and effective SSL approach in computer vision. Our approach utilises the proxy task of reconstruction, training a model to reconstruct intentionally corrupted signals. By applying this process across a large corpus of data, the model learns the salient representations necessary for accurate signal reconstruction. Similar to MIM, MSM involves several key design considerations, including the masking strategy $S_{\text{m}}$, masking ratio $R_{\text{m}}$, and model design. While the choice of model design is often influenced by external factors such as computational and memory requirements, $S_{\text{m}}$ and $R_{\text{m}}$ have a more explicit impact on the quality of pre-training [29].
We introduce several masking strategies for MSM, including random zero-masking (strat. A), random block zero-masking (strat. B), random noise-masking (strat. C), and block noise-masking (strat. D). Random masking involves randomly modifying I/Q samples with mask values (i.e., either zero or noise), while block masking randomly masks a continuous slice of the I/Q sequence. The extent of masking is governed by $R_{\text{m}}$, with higher values masking more of the signal and thus making the reconstruction task more challenging. As shown by MIM [29], random masking promotes the learning of local relationships between samples, while block masking promotes the learning of broader, global patterns from the data. Our signal masking process is formalised as
$s_{\text{m}}=\begin{cases}s\otimes 0_{\{X<R_{\text{m}}\}}&\text{if zero-masking},\\ s\oplus\hat{n}_{\{X<R_{\text{m}}\}}&\text{if noise-masking},\\ s&\text{if no masking},\end{cases}$ (1)
$\hat{n}\sim\mathcal{N}(\mu_{\text{train}},\sigma_{\text{train}}^{2}),\quad X\sim\mathcal{U}(0,1),\quad R_{\text{m}}\in[0,1],$ (2)
where $s$ and $s_{\text{m}}$ denote the original and masked I/Q sequences respectively, $\otimes$ and $\oplus$ denote sample-wise operations, and $\hat{n}$ denotes random noise sampled from the training distribution in the source domain.
III-C RF Datasets for Domain Adaption
We explore domain adaptation by applying MSM to I/Q signals from diverse RF domains, including telecommunications and radar signals. Our approach involves pre-training, without annotations, MAEs on four diverse datasets: RadioML [4], DeepRadar [10], RadarComm [9], and RadChar-SSL, which has been generated using our previous work [11]. We then perform few-shot transfer learning to fine-tune each pre-trained model on RadChar-nShot and evaluate performance on a test set (RadChar-Eval) derived from RadChar [11]. To ensure consistency across the pre-training datasets, we use only $10\%$ of RadioML, keeping dataset sizes comparable. An overview of dataset characteristics is provided in Table I.
TABLE I: RF Signal Classification Datasets
[…]
RadioML [4] Comm $255,590$ $1024$ $24$ $1$
DeepRadar [10] Radar $782,000$ $1024$ $23$ $0.01$
RadarComm [9] Mixed $125,361$ $128$ $6$ $0.1$
RadChar-SSL Radar $500,000$ $512$ $5$ $0.3$
RadChar-nShot Radar $205n$ $512$ $5$ $0.3$
RadChar-Eval Radar $100,000$ $512$ $5$ $0.3$
IV Experiments
IV-A Training Details
We perform pre-training, fine-tuning, and model evaluation on a single Nvidia Tesla A100 GPU. All models are trained with the Adam optimiser, where a constant learning rate of $0.001$ and $0.0001$ are used for self-supervised pre-training and fine-tuning, respectively. For pre-training, we train each model for $100$ epochs with early stopping based on validation loss with a $3$-step patience, using a fixed $70$-$20$-$10$ train-validation-test split for each dataset. For fine-tuning on limited frames, we train each model for the full $100$ epochs with the pre-trained encoder weights frozen for the first $10$ epochs. To maintain even class distribution, no validation split is considered during fine-tuning. We establish baseline few-shot performance for each model using the same model configuration as that used in fine-tuning, but without the pre-trained weight initialisation. A constant batch size of $128$ is used for pre-training, while a batch size of $8$ is used for both fine-tuning and baseline training. To improve generalisation and training stability, we standardise input signals with the mean and variance sampled from the training set of the respective RF domain. All models are evaluated on RadChar-Eval during testing.
IV-B Few-Shot Radar Signal Recognition
We establish a novel benchmark for few-shot RSR through the following experiments: (i) evaluating different MAEs; (ii) evaluating different masking strategies; (iii) evaluating the impact of SSL on test performance across different SNR settings; and (iv) evaluating the effectiveness of self-supervised pre-training on signals from different RF domains for few-shot radar signal classification. We evaluate the test performance of each linear probing classifier using classification accuracy and the macro F1 score. The test performance of each classifier is compared to its baseline, where the model is trained from scratch without self-supervised pre-training.
TABLE II: Summary of Few-shot Classification Performance on RadChar-Eval Dataset
Model Pre-Training 1-Shot Eval. (Best) 5-Shot Eval. (Best) 10-Shot Eval. (Best)
[…]
WaveNet RadChar-SSL A 0.6 $\mathbf{73.07}$ $\mathbf{70.63}$ A 0.8 $\mathbf{82.00}$ $\mathbf{80.26}$ A 0.8 $\mathbf{83.37}$ $\mathbf{81.64}$
DeepRadar D 0.2 $62.64$ $59.86$ A 0.8 $73.80$ $71.30$ A 0.8 $76.46$ $74.25$
RadarComm D 0.8 $67.09$ $64.43$ B 0.9 $75.79$ $73.53$ B 0.9 $77.83$ $75.70$
RadioML B 0.9 $64.84$ $61.62$ B 0.9 $77.40$ $75.24$ A 0.9 $78.66$ $76.57$
None - - $64.04$ $60.93$ $75.89$ $73.55$ $77.21$ $75.10$
MS-TCN RadChar-SSL A 0.7 $\mathbf{72.62}$ $\mathbf{70.13}$ A 0.9 $\mathbf{81.12}$ $\mathbf{79.27}$ A 0.9 $\mathbf{81.27}$ $\mathbf{79.40}$
DeepRadar C 0.6 $52.73$ $50.14$ C 0.6 $70.29$ $67.66$ C 0.7 $74.86$ $72.53$
RadarComm A 0.9 $68.40$ $65.72$ B 0.7 $76.13$ $73.89$ B 0.9 $78.25$ $76.14$
RadioML A 0.8 $64.76$ $61.95$ A 0.7 $75.36$ $72.96$ A 0.8 $77.89$ $75.75$
None - - $44.00$ $40.81$ $66.49$ $63.69$ $76.47$ $74.17$
ResNet RadChar-SSL A 0.7 $\mathbf{75.06}$ $\mathbf{72.32}$ A 0.7 $\mathbf{79.76}$ $\mathbf{77.76}$ A 0.8 $\mathbf{79.86}$ $\mathbf{77.73}$
DeepRadar A 0.4 $68.32$ $65.35$ B 0.2 $76.15$ $73.91$ B 0.2 $76.87$ $74.59$
RadarComm D 0.7 $70.03$ $66.98$ C 0.8 $77.21$ $75.00$ B 0.7 $78.05$ $75.82$
RadioML A 0.1 $74.30$ $71.81$ A 0.1 $77.35$ $75.16$ B 0.7 $78.08$ $75.96$
None - - $63.88$ $60.50$ $75.02$ $72.66$ $76.57$ $74.20$
Table II presents a summary of results for each pre-training configuration. The classification performance for models corresponding to the optimal $S_{\text{m}}$ and $R_{\text{m}}$ used during pre-training are shown. We observe that fine-tuned models that are pre-trained on data from the same domain (i.e., RadChar) yield the greatest improvement in test performance when compared to their respective baselines. This performance gain is most substantial in the $1$-shot configuration, which is the most challenging with the fewest labels available for fine-tuning. The improvement diminishes as the number of annotated frames used in fine-tuning increases, as reflected by a $17.5\%$, $6.3\%$, and $4.3\%$ increase in $1$-shot, $5$-shot, and $10$-shot performance for ResNet, respectively. Separately, we observe that MS-TCN’s baseline performance is relatively low in the $1$-shot setting, likely due to its larger model size ($33.6$ million parameters). This model requires more fine-tuning samples (Table II) to perform well compared to ResNet ($2.7$ million parameters) and WaveNet ($1$ million parameters).
(a) RadChar-SSL
(b) DeepRadar
(c) RadarComm
(d) RadioML
Figure 2: Impact of masking strategies and masking ratios on $1$-shot classification performance of ResNet, pre-trained across different datasets (a)–(d), and evaluated on RadChar-Eval.
(a) RadChar-SSL
(b) DeepRadar
(c) RadarComm
(d) RadioML
Figure 3: 1-shot classification performance of different models on RadChar-Eval across varying SNR levels. The optimal masking strategy and masking ratio used during pre-training are indicated in parentheses.
The optimal masking strategy and its corresponding masking ratio differ for each model and pre-training configuration. For ResNet (Fig. 2), random zero-masking (strat. A) with lower masking ratios (below $0.4$) performs best when pre-training on RadioML and DeepRadar, while the same strategy with a higher masking ratio (above $0.7$) is more beneficial when pre-training on RadChar-SSL. Separately, random block noise-masking (strat. D) with a higher masking ratio (above $0.7$) performs best when pre-training on RadarComm. The effectiveness of SSL on test performance depends on the SNR. Notably, models tend to benefit the most from pre-training when evaluating in moderate to high SNR levels (Fig. 3). We hypothesise that signal features (e.g., number of pulses, pulse width) are more prominent at higher SNRs, making reconstruction easier, whereas random noise dominates at lower SNRs, hindering reconstruction. In the most demanding fine-tuning configuration (i.e., $1$-shot), ResNet shows a performance boost ranging from $11.3\%$ to $25.7\%$ as SNR increases from $-20$ to $20$ $\mathrm{dB}$ when pre-trained on RadChar-SSL, whereas WaveNet’s improvement is more modest ($3.5\%$ to $6.4\%$) when compared with their respective baselines without SSL. WaveNet employs dilated convolutions, which provide a larger receptive field compared to ResNet. Consequently, WaveNet may benefit less from MSM, as the masking process may not present enough of a challenge for reconstruction. This is reflected in WaveNet’s preference for higher masking ratios (Table II).
While self-supervised pre-training on out-of-domain data generally yields varying degrees of performance improvement across different few-shot fine-tuning configurations (Table II), the effectiveness of RF domain adaptation depends on several key factors related to the extent of the domain gap between datasets. We hypothesise that differences in sampling rates (i.e., temporal resolution) between datasets affect how models learn the temporal relationships in the I/Q sequence. We suspect that pre-training on signals with higher temporal resolution ($t_{\text{res}}$) might lead to the model learning features that are less relevant when fine-tuned on signals with lower sampling rates, and vice versa. While resampling the source domain to match the target domain is possible, it reduces signal information and introduces aliasing, especially when the original signal has limited samples, such as in RadarComm (Table I). Conversely, upsampling requires interpolation, which may not accurately capture the characteristics of the original signal, potentially leading to bias and overfitting. Furthermore, differences in domain-specific characteristics and the number of signal classes ($N_{\text{cls}}$) in the source and target domains (Table I) can have a significant impact on fine-tuning performance. This is highlighted by comparing in-domain results in Fig. 3(a) with out-of-domain results in Fig. 3(b)–(d). Performance improvements from SSL are most notable for RadChar-SSL and RadarComm, where the number of signal classes ($6$ and $5$) and temporal resolution ($0.1$ and $0.3$) are most similar (Table I). In contrast, DeepRadar and RadioML contain substantially more RF classes than RadChar. Additionally, the signal features from DeepRadar and RadioML deviate significantly from the target domain. For example, the temporally smooth telecommunications signals from RadioML and the challenging low probability of intercept (LPI) radar classes (e.g., Px codes) in DeepRadar do not appear in the target domain, which may lead to over-regularisation and reduced effectiveness of fine-tuning. While RF domain adaptation is promising for few-shot RSR, ensuring that the signal characteristics in the source domain closely match those in the target domain is crucial for achieving effective and robust adaptation.
IV-C Ablation Study
We examine the impact of various design considerations for few-shot RSR. As discussed in Section IV-B, the choice of $S_{\text{m}}$ and $R_{\text{m}}$ influences SSL performance and, in turn, fine-tuning and test performance. Although no particular masking strategy consistently outperformed others across our experiments, we found that the masking ratio played a more important role in determining test performance (Table II). While a small batch size coupled with a low learning rate generally benefits fine-tuning, we observed that freezing the model weights for a few epochs provided a slight improvement (less than $5\%$) in test performance. This effect was consistent for models pre-trained on both in-domain and out-of-domain data. We also explored $\ell_{2}$ loss for pre-training, but we observed no meaningful improvements over $\ell_{1}$ loss in our experiments.
V Conclusion
In this paper, we introduced masked signal modelling as an effective self-supervised pre-training method for few-shot radar signal recognition. We also demonstrated the viability of RF domain adaptation for enhancing signal classification performance when no target domain data was used for pre-training. Our results show that by optimally designing the masking method during pre-training, fine-tuned models can achieve significant performance improvements, particularly in moderate to high SNR settings. This is demonstrated by the simple ResNet-based MAE, which achieved a boost in classification accuracy of $17.5\%$ when pre-trained on in-domain signals (i.e., RadChar-SSL) and $16.31\%$ when pre-trained on out-of-domain signals (i.e., RadioML), compared to its $1$-shot baseline without SSL. In future work, additional domains and downstream tasks (e.g., pulse activity segmentation) will be explored to assess the practical utility of SSL and RF domain adaptation for few-shot radar signal recognition.